### v2

@@ Abstract
The identification of art authors is crucial in areas like cultural heritage protection, art market analysis, and historical research. With the advancement of deep learning, Convolutional Neural Networks (CNNs) and Transformer models have become key tools for image classification. While CNNs excel in local feature extraction, they struggle with global context, and Transformers are strong in capturing global dependencies but weak in fine-grained local details. To address these challenges, this paper proposes a fusion model combining CNNs and Transformers for art author identification. The model first extracts local features using CNNs, then captures global context with a Transformer, followed by a feature fusion mechanism to enhance classification accuracy. Experiments on Chinese and oil painting datasets show the fusion model outperforms individual CNN and Transformer models, improving classification accuracy by 9.7% and 7.1%, respectively, and increasing F1 scores by 0.06 and 0.05. The results demonstrate the model’s effectiveness and potential for future improvements, such as multimodal integration and architecture optimization.
@@ 1 Introduction
Art identification, particularly in terms of author classification, is a critical research task in computer vision, with widespread applications in cultural heritage preservation, art market regulation, and historical research. With the rapid development of the internet, the digitization of artworks has enabled a large number of art pieces to be displayed and traded online. Consequently, automatic classification and identification of these artworks in terms of author have become key issues of interest in both industry and academia.
Author identification not only aids museums and galleries in accurately categorizing artworks but also provides technical support for the regulated management of the art market. With intelligent identification systems, users can efficiently filter artworks that meet their needs from a large collection. Moreover, the classification of art authors offers new perspectives for art scholarship, especially in areas like the authentication of artworks and the identification of creative periods, which have significant practical implications.
In recent years, deep learning techniques have made remarkable progress in image classification tasks, and Convolutional Neural Networks (CNNs) have become the mainstream approach for art author classification. CNNs automatically extract meaningful features from images through multi-layer convolution operations, overcoming the limitations and instability faced by traditional hand-crafted feature extraction methods (e.g., SIFT, HOG). The emergence of deeper networks, such as ResNet and Inception, has significantly improved image identification accuracy.
@@
    The model is tested on datasets from both Chinese painting and oil painting domains, enabling it to handle diverse art author classification tasks.
@@ 5 Conclusion
This study proposes an innovative fusion model that combines Convolutional Neural Networks (CNN) and Transformers to address the challenges of art author identification. CNNs, known for their exceptional ability to capture local features such as textures, brushstrokes, and color distributions, are used as the initial step to process the artwork images. Transformers, on the other hand, excel at modeling long-range dependencies and global relationships, which are crucial for understanding the overall composition and style of an artwork. By integrating these two powerful models, the proposed fusion approach leverages the strengths of both CNNs and Transformers, allowing the model to effectively extract detailed local features while also capturing the broader contextual information.

### v3

@@ Abstract
The identification of artwork is crucial in areas like cultural heritage protection, art market analysis, and historical research. With the advancement of deep learning, Convolutional Neural Networks (CNNs) and Transformer models have become key tools for image classification. While CNNs excel in local feature extraction, they struggle with global context, and Transformers are strong in capturing global dependencies but weak in fine-grained local details. To address these challenges, this paper proposes a fusion model combining CNNs and Transformers for artwork identification. The model first extracts local features using CNNs, then captures global context with a Transformer, followed by a feature fusion mechanism to enhance classification accuracy. Experiments on Chinese and oil painting datasets show the fusion model outperforms individual CNN and Transformer models, improving classification accuracy by 9.7% and 7.1%, respectively, and increasing F1 scores by 0.06 and 0.05. The results demonstrate the model’s effectiveness and potential for future improvements, such as multimodal integration and architecture optimization.
@@ 1 Introduction
Artwork identification, particularly in terms of author classification, is a critical research task in computer vision, with widespread applications in cultural heritage preservation, art market regulation, and historical research. With the rapid development of the internet, the digitization of artworks has enabled a large number of art pieces to be displayed and traded online. Consequently, automatic classification and identification of these artworks in terms of author have become key issues of interest in both industry and academia.
Artwork identification not only aids museums and galleries in accurately categorizing artworks but also provides technical support for the regulated management of the art market. With intelligent identification systems, users can efficiently filter artworks that meet their needs from a large collection. Moreover, the classification of artwork offers new perspectives for art scholarship, especially in areas like the authentication of artworks and the identification of creative periods, which have significant practical implications.
In recent years, deep learning techniques have made remarkable progress in image classification tasks, and Convolutional Neural Networks (CNNs) have become the mainstream approach for artwork classification. CNNs automatically extract meaningful features from images through multi-layer convolution operations, overcoming the limitations and instability faced by traditional hand-crafted feature extraction methods (e.g., SIFT, HOG). The emergence of deeper networks, such as ResNet and Inception, has significantly improved image identification accuracy.
@@
    The model is tested on datasets from both Chinese painting and oil painting domains, enabling it to handle diverse artwork classification tasks.
@@ 5 Conclusion
This study proposes an innovative fusion model that combines Convolutional Neural Networks (CNN) and Transformers to address the challenges of artwork identification. CNNs, known for their exceptional ability to capture local features such as textures, brushstrokes, and color distributions, are used as the initial step to process the artwork images. Transformers, on the other hand, excel at modeling long-range dependencies and global relationships, which are crucial for understanding the overall composition and style of an artwork. By integrating these two powerful models, the proposed fusion approach leverages the strengths of both CNNs and Transformers, allowing the model to effectively extract detailed local features while also capturing the broader contextual information.